### v3

# A Survey and Benchmark of Automatic Surface Reconstruction from Point Clouds

Abstract
We present a comprehensive survey and benchmark of both traditional and learning-based methods for surface reconstruction from point clouds. This task is particularly challenging for real-world acquisitions due to factors like noise, outliers, non-uniform sampling, and missing data. Traditional approaches often simplify the problem by imposing handcrafted priors on either the input point clouds or the resulting surface, a process that can necessitate tedious hyperparameter tuning. Conversely, deep learning models have the capability to directly learn the properties of input point clouds and desired surfaces from data.
We study the influence of these handcrafted and learned priors on the precision and robustness of surface reconstruction techniques. We evaluate various time-tested and contemporary methods in a standardized manner. When both trained and evaluated on point clouds with identical characteristics, the learning-based models consistently produce superior surfaces compared to their traditional counterparts—even in scenarios involving novel shape categories. However, traditional methods demonstrate greater resilience to the diverse array of point cloud anomalies commonly found in real-world 3D acquisitions. For the benefit of the research community, we make our code and datasets available, inviting further enhancements to learning-based surface reconstruction. This can be accessed at https://github.com/raphaelsulzer/dsr-benchmark.
Index Terms:
surface reconstruction, point clouds, deep learning, mesh generation, survey, benchmark1 Introduction
Modern three-dimensional (D) acquisition technology, such as range scanning or multi-view stereo (MVS) allows us to capture detailed geometric information of real scenes in the form of point clouds. However, sets of D points are usually not sufficient for modeling complex physical processes, such as fluid dynamics or environmental phenomena. A variety of applications in science and engineering require a continuous surface representation of objects or scenes. As a result, transforming point clouds into surface models—a process known as surface reconstruction—is a central problem in digital geometry processing. In this paper, we offer a comprehensive survey and benchmark of traditional and learning-based methods designed to address this problem.
Without prior information about the sought surface, the task of surface reconstruction from point clouds becomes ill-posed: there are infinitely many surfaces that can potentially fit the given points. As illustrated in Fig. 1, acquisition defects such as non-uniform sampling, noise, outliers, or missing data further complicate the reconstruction of a geometrically and topologically accurate surface [1].
Traditional surface reconstruction techniques often simplify this problem with handcrafted priors on the input cloud: point density, noise and outlier levels, or the output surface: smoothness, topological properties, or semantics. However, adapting priors to different acquisitions often requires tedious hyperparameter tuning. In contrast, contemporary deep learning methods learn point cloud defects and shape properties directly from training data. In consequence, they can reconstruct accurate surfaces without manual parameter tuning. However, neural surface reconstruction (NSR) methods have mainly been tested on artificial data sets with a limited number of object categories. Such datasets may not be representative of the endless variety of shapes in real-world scenarios. Moreover, NSR methods are typically applied on uniformly sampled point clouds. These do not reflect the irregularities of real-world acquisitions such as non-uniformity, missing data from occlusions, or transparent regions. The ability of NSR models to reconstruct shapes in unseen shape categories and from point clouds with unseen defects has rarely been rigorously and systematically evaluated.
In this paper, we propose a series of experiments designed to benchmark surface reconstruction algorithms from point clouds. We use a diverse range of publicly available shape datasets with objects of varying complexities and for which we know the true surface. To emulate real-world scenarios, we synthetically scan the objects to generate point clouds with realistic characteristics. Working with objects with known surfaces allows us to measure the geometric and topological quality of the reconstructed surfaces. We also consider point clouds from real acquisitions for which the true surface is typically unknown. By comparing novel learning-based algorithms to traditional test-of-time methods, we specifically study the impact of learned v.s. handcrafted priors in the reconstruction process. Throughout our study, the generalizability of these methods remains a focal point. Our main contributions are as follows:
We provide a comprehensive literature review of surface reconstruction methods from point clouds, tracing developments over three decades up to the latest learning-based approaches. We contrast popular test-of-time traditional models with novel NSR models. Additionally, we discuss their relation to the new emerging field of rendering-based surface reconstruction.
We benchmark both traditional and learning-based methods under consistent conditions. We perform several experiments to assess their geometric precision, topological consistency, and robustness using openly available shape datasets and point clouds from both synthetic and real scans.
2 Related work
In this section, we provide an overview of the existing surveys and benchmark studies of NSR methods.
2.1 Surveys
Although several works survey the broad field of surface reconstruction from point clouds [2, 1, 3, 4, 5, 6], the majority predate the rise of learning-based reconstruction. Surface reconstruction methods are often grouped as interpolation-based [2] or approximation-based techniques [1]. Interpolation methods “connect” points of the input point cloud, usually by linear interpolation between pairs of points. Approximation methods use smooth functions to represent the point cloud globally or locally [7]. In our review, we consider both interpolating and approximating methods and focus on novel ideas in learning-based surface reconstruction. While many reconstruction methods can be distinguished by their surface priors [1, 5], we argue that several successful approaches combine different priors, complicating classification solely based on priors. We thus organize the methods into two groups: surface-based and volume-based approaches. This breakdown mirrors the primary mathematical surface representations: parametric [8] and implicit [9]. To the best of our knowledge, only two surveys include recent learning-based methods [5, 10]. The work of You et al. [3] predates important developments for learning-based methods, such as the incorporation of local information [11, 12, 13, 14, 15, 16, 17].
2.2 Benchmarks
Historically, very few benchmarks for surface reconstruction from point clouds have been proposed. Numerous methods use bespoke datasets to evaluate their approach, typically by uniformly sampling point clouds from the ground truth surface from existing shape collections [18, 11, 12, 13, 16, 17]. However, the characteristics of these sampled point clouds often differ between studies, complicating the comparison of their results. Moreover, these point clouds typically lack common defects seen in real acquisitions, such as missing data from occlusion.
Two notable exceptions are the benchmarks of Berger et al. [19] and Huang et al. [5]. Both teams develop synthetic range scanning procedures to produce point clouds with realistic artifacts such as noise, non-uniformity, and misaligned scans. They also use shapes with non-trivial topology and details of various sizes. While providing interesting results, Berger et al.’s benchmark predates learning-based surface reconstruction and only considers traditional approximation techniques. Huang et al.’s benchmark considers learning-based methods, but does not specifically study their generalization capabilities.
In our benchmark, we use the synthetic range scanning procedure of Berger et al. and their test shapes as they provide a realistic challenge for both learning-based and traditional algorithms. We also implement our own synthetic scanning procedure to produce MVS-like point clouds. We use synthetic scanning on existing large shape datasets to create training datasets with true surfaces and point clouds with realistic characteristics (see Fig. 5). Furthermore, we specifically study the generalization capabilities of methods by using train and test sets with different characteristics in terms of shape categories and acquisition defects.
MVS benchmarks. The generation of point clouds from 2D information such as overlapping images is an integral problem of surface reconstruction. There exists a variety of benchmarks using data captured in a laboratory environment [20, 21] or in the wild [22, 23, 24]. These benchmarks often use low-quality image acquisitions as input, and use a higher-quality acquisition, e.g., from LiDAR scans, to produce precise and dense point clouds that serve as reference. However, even with advanced sensors, producing a complete and precise point cloud remains difficult. As a result, the evaluation of the reconstructed surfaces can be flawed. A common workaround is to restrict the evaluation to specific “trustworthy” areas [21, 23]. However, in contrast to true and complete surfaces, this approach prevents the computation of topological metrics such as component count, or differential metrics such as surface normals. Furthermore, many learning-based methods require a large dataset of closed reference surfaces and not just a small set of precise point clouds.
3 Representations, properties, and reconstruction of surfaces
In this section, we first define the notion of surface and its mathematical and digital representations. We then discuss the most critical properties of reconstructed surfaces. Finally, we propose a grouping of surface reconstruction algorithms featured in our survey based on the mathematical representations of surfaces: parametric or implicit.
Throughout this article, we denote by the input point cloud, the reconstructed surface, and the real surface from which is sampled.
3.1 Representations
A surface is defined as an orientable, continuous 2-manifold within , which may or may not have boundaries [7, 25, 26]. These properties hold significance for surface visualization and processing, which we will elaborate on in subsequent sections. From a mathematical perspective, surface representations fall primarily into two categories: parametric and implicit.
Parametric surfaces. These surfaces are represented by a continuous and bijective function that maps a parameter domain (e.g., ) to a 3D surface . However, it may be impractical to devise a single function to parameterize complex surfaces. Therefore, the parameter domain is usually divided into smaller sub-regions, each equipped with its own function [7]. Typically, is partitioned into triangles. A set of triangles approximating can be efficiently stored and processed as a triangle surface mesh , where denotes the vertices of the triangles, their edges and their facets.
Implicit surfaces. These surfaces are defined by the level-set of a scalar-valued function :
The most common choice for are signed distance function (SDF) or occupancy function (OF). An SDF gives the distance from a D point x to the closest part of the surface. Points within the interior of the surface are assigned negative values, while those outside are given positive values. An OF typically has a value of inside the surface and outside. The -level-set of represent the surface with for SDFs and for OFs. Similar to the parametric case, the domains of implicit functions are often split into sub-regions, such as voxels [11], OcTree-nodes [27], or tetrahedra [28]. In practice, once computed, implicit surfaces are often converted into a parametric mesh, which is easier to handle for downstream applications.
3.2 Properties
We want the reconstructed surface to be close in terms of geometry and topology to the real surface from which the point cloud is sampled. Real surfaces exhibit certain intrinsic properties, such as watertightness, manifoldness, and orientability, which are crucial for a variety of downstream applications. Therefore, it is essential that the reconstructed surface also possess these properties. In this section, we define these properties in the context of mesh representations. For a visual interpretation of these concepts, the reader is directed to Fig. 2.
[…]
Manifoldness: Real and geometric surfaces are 2-manifold if and only if any of its point has a local neighborhood that is homeomorphic to an open subset of the Euclidean plane. In other words, every small enough region of the surface, irrespective of its curvature, can be continuously deformed into a 2D plane. A mesh is manifold if it has the following attributes:
[…]
Vertex-manifold: For each vertex in , the set of facets that share this vertex forms a topological half-disk. This implies that facets that share a common vertex are organized in an open or closed fan. In other words, each vertex is part of at least two edges and each edge is part of two facets.
[…]
The watertightness property is crucial for physical simulations such as fluid dynamics. Both manifoldness and orientability are prerequisites for certain mesh data structures, such as the widely used half-edge structure [29, 30]. These structures ensure consistent and predictable behavior during geometric transformations, topological modifications, and various mesh optimization and query operations, and many mesh processing algorithms require a manifold mesh. Lastly, intersection-free and orientable surfaces lead to a well-defined notion of inside and outside, which is important for mesh visualization and various geometric operations and metrics.
3.3 Reconstruction
Surface reconstruction from point clouds consists in creating a continuous surface approximating the original surface from which the 3D points were sampled. In our survey, we group the methods for surface reconstruction from point clouds into two groups: surface-based and volume-based.
Surface-based methods. These approaches focus on finding one or several parametric surfaces approximating the input point cloud , either in the form of triangles or two-dimensional (D) patches, or by deforming parameterized enclosing envelops such as meshed spheres. The main limitation of surface-based methods with a single function f is that the topology of the parameter domain must be equivalent to the topology of the true surface , which is usually unknown. Conversely, the main challenge for surface-based methods that define with multiple functions for different regions is to guarantee consistent transitions between subsurfaces. In particular, it is difficult to keep the final surface intersection-free, manifold, and watertight.
Volume-based methods. These methods segment a subset of into interior (inside) and exterior (outside) subspaces. The surface is implicitly defined as the interface between the two subspaces. Most, but not all algorithms in this class formulate the problem as finding an implicit function indicating where a point in space is full (inside) or empty (outside). Surfaces reconstructed with volume-based methods are guaranteed to be watertight and intersection-free, but not necessarily manifold [2].
qtree/.style=for tree=parent anchor=south, child anchor=north,align=center,inner sep=0pt,l=1.2cm
[…]
Mesh extraction. Surface-based methods typically directly yield a mesh, e.g., by triangulating . In contrast, volume-based methods usually require an additional processing step. If the implicit field is discretized with tetrahedra using 3D Delaunay tetrahedralisation (DT), a mesh can then be built from all triangles that are adjacent to one inside-tetrahedra and one outside-tetrahedra. Another option is the algorithm of Boissonnat and Oudot [59] that iteratively samples the implicit function by drawing lines traversing the volume and building a triangle mesh from their intersection with the underlying surface . One of the most popular methods for extracting a mesh from an implicit field is Marching Cubes [60], which discretizes the implicit function into voxels, constructs triangles inside each voxel that have at least one inside and one outside vertex, and extracts a triangulation as the union of all triangles. Recently, mesh extraction has also been addressed with deep learning methods [61, 62, 63, 64]. Neural meshing [64] specifically addresses the case where an implicit function is encoded in the weights of a neural network and aims to extract meshes with fewer triangles compared to Marching Cubes. A novel interesting approach in the direction of mesh extraction is the use of Voronoi diagrams to extract concise polygon meshes from an implicit function [61].
Robustness. When the point cloud is sampled densely and without noise, both surface-based and volume-based methods can provide theoretical guarantees about the topology and geometry of the reconstructed surface [2]. However, in this paper, we focus on the robustness of various methods to defect-laden input point clouds from D scanning.
Input. The characteristics of the input point cloud mainly depend on the sensor used for their acquisition.
LiDAR point clouds typically offer high precision and accuracy without a significant amount of noise or outliers due to the robustness of the active sensor to challenging conditions. However, LiDAR point clouds commonly suffer from occlusion due to the sensor’s limited field of view.
Image-based point clouds are produced using a technique called multi-view stereo (MVS). Their quality heavily depends on the quality and quantity of input images. Pixels from overlapping images are matched which allows to compute their 3D location relative to the camera position. MVS can produce dense point clouds, especially in areas with abundant texture and features. However, poor lighting conditions or lack of texture can result in less accurate point clouds, and variations in image scales to strongly varying point densities [23].
4 Survey
In this section, we review important surface- and volume-based surface reconstruction methods and discuss their robustness against different types of point cloud defects. We also explore the numerous links between learning-based and traditional methods. Finally, we discuss recently emerging rendering-based methods for surface reconstruction.
4.1 Surface-based reconstruction
Surface-based approaches directly produce a surface, typically as a mesh, by connecting some of the points of or deforming a reference mesh.
4.1.1 Surface-based interpolation
Most traditional surface-based approaches linearly interpolate between all or a selection of 3D points of . A critical aspect lies in how these points are connected into triangles, leading to a variety of algorithms.
Advancing-front. Computing the DT of can be done by forming triangles for all triplets of points respecting the empty ball property, i.e., such that no other points lie within their circumsphere, which is the smallest sphere passing through three non-co-linear points. The Ball-pivoting algorithm [31] employs a greedy strategy to find such triplets and consists of several steps. First, it selects a seed triplet of points with the empty ball property such that their circumsphere has a radius close to a defined value related to the density of . This triplet defines the first triangle of the mesh, and its circumsphere forms a ball used for the rest of the algorithm. This ball is then pivoted around each edge of the triangle until it touches a new point, forming a new triplet, and hence a new triangle. Once all possible edges have been processed, the algorithm starts over with a new seed triangle until all points of have been considered.
A similar approach to the Ball-pivoting algorithm, avoiding the fixed ball size, is the regular interpolant method [32]. It also starts by triangulating a seed triplet of points that adheres to the empty ball property. Then, for each existing edge a new triangle is added to the triangulation. The point is chosen such that the new triangle has the smallest radius circumsphere for all .
A refinement of the Ball-pivoting algorithm [33] also use a ball of varying size, and additionally proposes a parallel implementation. Both follow-up works have enhanced the Ball-pivoting algorithm’s robustness against nonuniform sampling. However, these methods remain sensitive to point cloud defects such as noise or large occlusions.
Selection-based. Similar to advancing-front techniques, some learning-based methods [35, 34] also iteratively build the triangulation from an initial set of candidate triangles. PointTriNet [34] computes the -nearest neighbor graph of and creates seed triangles. Then, a first network selects some of these triangles based on their probability of being part of the final surface, which is obtained by considering their neighbouring points and already-formed triangles. Finally, a second network proposes new candidate triangles formed by points adjacent to selected triangles. The new candidates are processed by the first network and the algorithm continues for a user-defined number of iterations. As the loss function is the Chamfer distance between the input points and the reconstructed surface, this method can be trained without ground-truth meshes.
The intrinsic-extrinsic ratio (IER) between two points is the ratio of the Euclidean distance and the geodesic distance along the true surface. An IER value near for a pair of points suggests that the edge connecting them is likely to be part of the surface. IER-meshing [35] starts with a large set of seed triangles generated from the -nearest neighbor graph of the point cloud. They then estimate the IER for the vertices of these triangles using an multilayer perceptron (MLP) supervised from a ground-truth mesh. Triangles whose vertices exhibit IER values close to are then selected for the mesh, as this indicates a high probability of their edges being part of the actual surface. While they are robust against minor noise to the input point clouds, the aforementioned methods typically result in reconstructed surfaces that are neither manifold nor watertight.
Projection-based. A notable example of projection-based methods is the algorithm by Boissonnat [36], which may be the earliest surface reconstruction algorithm [2]. This method starts by computing the tangent plane of each point, projecting their neighboring points onto these planes, computing a 2D Delaunay triangulation of the projected 2D point clouds, and merging these local reconstructions.
The use of tangent planes is difficult in areas with high curvature or for reconstructing thin structures. DSE-meshing [16] addresses these challenges by employing logarithmic maps as an alternative. Logarithmic maps offer local surface parametrizations around a point based on the geodesics from that point, providing a more adaptable approach for complex geometries. This method uses an MLP to obtain approximate logarithmic maps for each point. Finally, neighboring maps are aligned and triangulated. This approach leads to fewer non-manifold edges compared to methods that process triangles independently. However, the surface may still not be watertight.
4.1.2 Surface-based approximation
Patch fitting. AtlasNet [37], one of the pioneering learning-based methods in surface reconstruction, employs an original approach. It starts with a triangulated 2D patch and uses a neural network to adjust its vertices’ positions. This network is trained by minimizing the distance between the deformed surface and the point cloud, eliminating the need for actual mesh data for training purposes. Similar to interpolating methods, AtlasNet does not inherently ensure that the reconstructed surface is both watertight and manifold.
Mesh deformation. A variation of the same idea is the mesh-fitting approach, which starts with an intial mesh based on some of the input points [39] or a low-resolution Poisson [27] reconstruction [40], and adjusts the position of the vertices to fit the point cloud. Such methods demonstrate robustness in scenarios with missing data, but they require meticulous parameter tuning to effectively manage noise or outliers. A notable limitation is the fixed topology of the initial mesh throughout the reconstruction process. Consequently, if the initial mesh does not accurately reflect the true topology of the target surface, this discrepancy cannot be corrected later in the process. For example, if the initial mesh is convex and the desired surface includes holes, it can not be reconstructed with consistent orientation. A significant advantage of mesh deformation methods compared to other surface-based approaches is that they guarantee a watertight surface.
4.2 Volume-based reconstruction
Instead of directly generating a surface mesh, volume-based approaches consider discrete or continuous 3D volumes from which a mesh can be extracted. These methods offer a different perspective on the reconstruction process, initially emphasizing volume over surface geometry.
4.2.1 Volume-based interpolation
[…]
-shapes. One straightforward way to extract such a subcomplex from a DT involves two key steps: discarding all tetrahedra whose circumspheres are larger than a predetermined radius , and then retaining only the boundary triangles, i.e., triangles between tetrahedra that are discarded and tetrahedra that are kept, resulting in the point cloud’s -shape [41]. Like the Ball Pivoting algorithm, the selection of depends on the density of points. For samplings that are error-free and dense, -shapes, along with other interpolation methods [2, 66, 36], can offer provable guarantees of topological accuracy of the reconstructed surface [2]. A recent approach constructs a DT that is greedily refined and carved with an empty ball of size to reconstruct a watertight 2-manifold surface [42]. The surface strictly encloses the input point set by a positive user-defined offset.
[…]
4.2.2 Volume-based approximation
Implicit functions. Implicit functions form the cornerstone of many popular volume-based surface approximation techniques. The seminal work by Hoppe et al. [25] pioneered this approach by defining an SDF based on the local tangent planes at each point. However, the computation of these tangent planes proved costly and susceptible to noise, as well as to variations in point density [51]. Over the years, implicit function methods have been refined and improved to address these challenges.
Arguably the most well-known method for surface reconstruction, Poisson surface reconstruction was introduced in 2006 by Kazhdan et al. [50]. The method uses an indicator function and a level set that defines the reconstructed surface . The main idea of the approach is that the Laplacian of should equate to the divergence of a vector field defined by the oriented normals of the points of :
The reconstruction process starts by building an octree from and defines a system of hierarchical functions for each cell of the tree. The function coefficients are computed by solving a sparse linear system, making the method time and memory-efficient.
While Poisson reconstruction yields a closed and smooth surface, it tends to erase small details and structures. Screened Poisson Surface Reconstruction (SPSR) [27] addresses this problem by constraining the implicit function, and thus reconstructed surface, to pass through or near all points of . Moreover, this refined variant adds Neumann boundary conditions [72], which allow the reconstructed surface to intersect the boundary of the domain on which is defined. This was further refined by adding Dirichlet boundary conditions around a tight envelope enclosing the point cloud and leading to better reconstructions in areas of missing data [52]. Poisson methods generate watertight meshes and are robust against a wide array of acquisition defects of moderate amplitude. However, a notable prerequisite is the need for well-oriented normals, which can be a significant challenge in real-world data acquisition.
Rather than being characterized by its Laplacian and boundary conditions, the implicit function can be represented as a neural field with a neural network [53, 18, 73]. This broad category of methods is further subdivided into three distinct subgroups based on the nature of the network employed: set-based, grid-based, or neighborhood-based. Each subgroup represents a unique approach to leveraging neural networks for the representation of implicit functions.
Set-based neural implicit functions. This first class of approaches consider the input point cloud as an unordered set of points and compute a global embedding for its global structure. Occupancy Networks (ONet) [53] defines as a fully-connected network (FCN) conditioned by the input point cloud . This network is trained to predict whether a set of preset points lie inside or outside the surface . To accurately train the network, true watertight surfaces are required for reference, such as those found in some CAD-based object databases.
DeepSDF [18] takes a different approach by conditioning on a latent shape code optimized for a given input point cloud during inference. This optimization is facilitated by an auto-decoder network trained on a collection of shapes. However, the optimization process requires accurate signed distance values for the input point cloud, which can be challenging for complex or ambiguously oriented shapes.
The Local Implicit Grid (LIG) [13] and DeepLS [12] approaches split the input point cloud into overlapping subregions and process them independently with weight-sharing networks. A benefit of this approach is that similar areas belonging to different objects, such as flat surfaces on tabletops or floors, are processed uniformly, thus reducing the risk of the network overfitting to specific training scenes. However, as DeepSDF, these techniques also require true values of the implicit function for the shape during inference. Furthermore, the dimension of the subregions is a critical parameter that requires careful calibration to ensure successful reconstruction.
Implicit Geometric Regularisation (IGR) [54] proposes a different approach in which a FCN is optimized during inference for each test shape such that the signed distance’s -level set goes through the points of and its Laplacian follows the points normal. As this approach is optimized for each shape independently, it does not require true meshes for supervision, but it also cannot learn shape priors.
Grid-based NIFs. As they rely on a single point cloud embedding to define the implicit function , set-based NIF methods tend to discard local point cloud information, which often results in oversmoothing [11]. To address this issue, the 3D space can be discretized into a voxel grid which can be used to learn local features.. Then, D or D U-Nets [74, 75] process the resulting maps to capture local and global shape information and, in turn, connect a fully-connected network predicting occupancy. Convolutional Occupancy Networks (ConvONet) [11] exemplifies this approach and also proposes a sliding-window strategy to scale to larger scans. However, since the content of the windows must be consistent with the scale of training shapes, this approach can be computationally expensive for large scenes.
Using a similar encoder architecture, Shape As Points (SAP) [55] merges neural implicit fields with Poisson reconstruction techniques. The process starts by densifying the input point cloud and estimating the point normals. This enhanced point cloud is then fed into a differentiable Poisson solver, which determines the occupancy for each cell of a regular grid (cf. the indicator function of Poisson). These predicted values are then compared to true occupancies with the L2 loss, leading to a fully differentiable pipeline where point encoding, densification, and normal estimation are all learned end-to-end. Shape As Points actually comes in two flavors: a purely optimisation-based Shape As Points (SAP∗), which does not use any training, and a learning-based setting, for which we retain the name SAP.
Neural Kernel Fields (NKF) [57] also combine learned features with techniques of the classical Poisson surface reconstruction. Input points are encoded into a feature grid, which is then used to compute local kernel basis functions for every point. Similar to PSR, the function coefficients are computed by solving a sparse linear system and using occupancy information from oriented normals. Recently, NKF has been updated to make the method more scalable and robust to noise, by using hierarchical features from multiple voxel grids at different resolutions and aligning the implicit function gradients with the normal direction of the input points [58, 56]. In contrast to the original Poisson Surface Reconstruction, the aforementioned methods allow to incorporate learned priors into the implicit function computation. Additionally, for SAP the input point cloud does not need to be equipped with oriented normals.
Neighborhood-based NIFs. The choice of the voxel size for 3D convolution-based methods is often critical: opting for small voxels increases computational and memory requirements, whereas coarse grids discard geometric details. Points2Surf (P2S) [14] proposes an alternative approach that predicts signed distance functions by encoding both the local neighborhood of each point (for unsigned distance prediction) and a downsampled representation of the entire shape (for orientation prediction). Leveraging -nearest neighbor sampling, this method offers better adaptability to variations in point density. However, it may result in extended processing times since the local neighborhoods of each point needs to be processed individually during the inference stage. Point Convolution for Surface Reconstruction (POCO) [17] proposes instead to use a continuous convolution backbone [76] to learn point features without spatial discretization. The occupancy of each point is then determined by an attention-based scheme.
4.3 Rendering-based reconstruction
Our paper specifically addresses surface reconstruction methods from 3D point clouds. However, a new emerging field is the reconstruction of surfaces directly from images using rendering techniques from computer graphics. Originally, so-called novel view synthesis methods [77, 78] were specifically designed for generating new views of an object or scene starting from a set of given images with their accurate pose. These methods present many advantages, including the ability to generate photo-realistic content, to deal with complex lighting and transparencies [79], and the fact that they do not rely on complex sensors. Recently, more and more rendering-based methods have been designed specifically for surface reconstruction [80, 81, 82]. They can be divided into the two categories, neural radiance fields and 3D Gaussian splatting and are closely related to the two corresponding, more general, principles from computer graphics, namely ray tracing and rasterization.
4.3.1 NeRF approaches
NeRF encode a radiance field in a neural network. The network maps from each 3D location and 2D viewing direction to a volume density and directional emitted color [77]. The color of any ray passing through the scene is rendered using principles from classical volume rendering [83]. Rendering rays from the given input images and poses allows to train the network weights, and to generate novel views from the trained network. Similar to the NIF methods that encode signed distance, or occupancy fields presented above, the volume density of neural radiance fields can be used to infer an object’s surface. However, the trivial solution of extracting the surface as a level-set of the volume density of a NeRF does not lead to a satisfying surface [80]. A common approach is to force the volume density field to change abruptly, indicating the presence of surface [81].
The benefit of ray tracing-based approaches is that the two-step approach of point cloud generation from images and subsequent surface reconstruction from the point clouds is solved in one end-to-end optimization. The downside is that the methods are often slow due to frequent querying of the radiance field during optimization.
4.3.2 3D Gaussian splatting approaches
A more efficient rendering technique compared to ray tracing is rasterization. 3D Gaussian Splatting [84] is the first method that makes use of a differentiable rasterization to generate novel views of an object or scene. The method takes a sparse point cloud, typically from SfM as input. The points are then used to initialize 3D Gaussians which are rendered using rasterization. The position and density of the 3D Gaussians is optimized using an image-based loss between input and rendered views.
The method is significantly faster for novel view synthesis compared to NeRF. However, for surface reconstruction, 3D Gaussians still need to be processed in a seperate step, e.g., using Poisson surface reconstruction [82], leading to a two step solution, similar to a traditional MVS pipeline.
5 Benchmark setup
In this section, we propose a series of experiments to benchmark surface reconstruction algorithms discussed in the previous section. We first describe our approach to generate realistic point clouds (Sec. 5.1) and detail the different datasets that we use (Sec. 5.2). We then define the experimental setup (Sec. 5.3), the methods evaluated (Sec. 5.4) and finally the benchmark’s metrics (Sec. 5.5).
5.1 Synthetic scanning for point cloud generation
In an ideal scenario, we would use real point cloud acquisitions and compare the reconstructed surfaces with the associated true surfaces. However, it is often impractical to obtain the true surfaces of real objects, as it requires high-density and error-free scans as well as substantial manual intervention. Many existing MVS benchmarks derive their point clouds from images, and the true surfaces are obtained with extensive acquisitions from multiple stationary Light Detection and Ranging (LiDAR) scans [22, 20, 21, 23, 24]. However, even these surfaces may be incomplete or not watertight due to occlusions. Additionally, these datasets are typically limited in scale, featuring only a few objects or scenes. As a result, while they are valuable for evaluation purposes, they are less suitable for training models. Instead, we propose to use synthetic surfaces from extensive shape collections for training, since they come with strong topological guarantees. In order to reproduce the defects commonly encountered during real scans, such as density variations and missing data from occlusion, we implement realistic artificial scanning procedures, illustrated in Fig. 6.
5.1.1 Synthetic MVS
We propose a procedure designed to replicate the typical characteristics of point clouds obtained from MVS techniques. While our approach doesn’t follow an MVS pipeline, it aims to produce point clouds with similar defects as those found in MVS-generated point clouds. However, it does not model failure cases from MVS methods, such as the lack of points in areas with difficult lighting conditions or without texture, or strong variations in density and noise levels.
We randomly position virtual sensors around a sphere enclosing the object, and shoot rays towards random points on the object’s circumsphere. A new 3D point is recorded at the first intersection of the ray with the object’s surface. Further intersections are ignored, leading to similar occlusion patterns as in real world acquisitions. We create two types of scans using this methodology: (i) sparse point clouds, consisting of approximately 3,000 points per object; (ii) dense point clouds, comprising around 10,000 points per object, with 10% of these points being outliers. Consistent with the approach suggested by Peng et al. [11], we apply Gaussian noise to each point position, characterized by a zero mean and a standard deviation of 0.5% of the bounding box diagonal. Both sparse and dense scans are conducted from 10 uniformly distributed sensor positions.
5.1.2 Synthetic range scanning
We use a modified version of the range scanning procedure from the surface reconstruction benchmark of Berger et al. [19]. We choose five distinct scanning settings, each presenting varying levels of difficulty, to scan each shape:
LR: Low resolution to replicates point clouds obtained from long-range scanning;
HR: High resolution with close to no defects;
HRN: Noisy acquisition by adding random jitter on the point positions;
HRO: Outliers by adding points uniformly sampled the bounding box of the object;
HRNO: Noise and outliers combines the last two defects.
See the supplementary material for the exact numbers and the original benchmark paper [19] for further details.
5.2 Datasets
To train and evaluate various surface reconstruction algorithms from point clouds, we use a mix of synthetic and real shapes. See the supplementary material for visual examples.
Training sets. To have enough shapes to train deep networks, we use the synthetic and closed shapes from ShapeNet [85] and ModelNet [86]. Point clouds are generated using our MVS procedure, and the true surfaces are processed with ManifoldPlus [87] to make them watertight. We generate point clouds with 3000 and 10,000 points for ScanNet, and 3000 points for ModelNet as they are typically simpler models. Following the standard practice in this field, we use Choy et al.’s [88] 13-class subset of ShapeNet and the corresponding train/val/test split. We use the training sets to parametrize non-learning-based methods, ensuring a fair comparison across different algorithms.
Evaluation. We evaluate the methods using five shapes from the benchmark of Berger et al. [19] with challenging features, such as intricate details and complex topology. These shapes present a higher reconstruction difficulty compared to those in ModelNet. For each shape, we generate point clouds: with 3000 and 10,000, and using synthetic MVS and range scanning procedures. For qualitative evaluation, we also consider real scans, including open surfaces. We select a range scan from Tanks and Temples [24] (Truck), and two MVS point clouds from DTU [21] (scan1) and from Middlebury [20] (Temple), each subsampled to 50,000 points.
5.3 Design of the experiments
| Experiment | Training set | Test set | ||
| Evaluation of methods with dataset-driven parameterization | ||||
| E1 In-distribution |
| ShapeNet (synthetic MVS) | ShapeNet (synthetic MVS) | |||
| E2 Out-of-distribution unseen point cloud characteristics: density, outliers |
| ShapeNet (synthetic MVS) | ShapeNet (synthetic MVS) | |||
| E3 Out-of-distribution unseen shape categories: less complex shapes |
| ShapeNet (synthetic MVS) | ModelNet (synthetic MVS) | |||
| E4 Out-of-distribution unseen shape categories: more complex shapes |
| ModelNet (synthetic MVS) | ShapeNet (synthetic MVS) | |||
| Evaluation of methods without dataset-driven parameterization | ||||
| E5 Shape variety with defects shapes with varying complexities, density, noise and outliers | – | – |
| Berger et al. (synthetic range scan) | ||||
| Evaluation of all methods on a common ground | ||||
| E6 Synthetic data with ground truth (quantitative evaluation) |
| ShapeNet (synthetic MVS) | Berger et al. (synthetic MVS) | |||
| E7 Real data without ground truth (qualitative evaluation) |
| ShapeNet (synthetic MVS) | Middlebury (MVS), DTU (MVS), T&T (range scan) | |||
We design a series of experiments (Tab. I) to thoroughly evaluate surface reconstruction models.
Evaluation of methods with dataset-driven parameterization. We benchmark both learning-based methods and traditional optimization-based methods that have parameters that can be set according to a given dataset and which can be evaluated on a complete dataset.
(E1) In-distribution. We trained models on the ShapeNet training set and test on its test set. All point clouds are generated using the same MVS scanning procedure, featuring 3000 points with Gaussian noise (zero mean, standard deviation of 0.5% of the bounding box diagonal). This experiment assesses the models’ ability to handle sparse point clouds, to complete missing data, and to eliminate noise.
(E2) Out-of-distribution - unseen point cloud characteristics. We test models trained in E1 on denser point clouds with 10,000 points with the same 0.5% Gaussian noise, and with the addition of 10% outliers.5 This tests the ability of the model to generalize to different point cloud characteristics.
(E4) Out-of-distribution - unseen shape categories: more complex. We train the learning methods on the simpler shapes from ModelNet and evaluate them on the test set of ShapeNet. Here, we assess whether learning methods can generalize from simple shapes to more complex ones, a difficult out-of-distribution setting.
Evaluation of methods without dataset-driven parameterization. As the tested neural optimization-based methods are significantly slower than learning-based and traditional optimization-based methods (several minutes v.s a few seconds for a single shape), we had to evaluate them separately, on a much smaller dataset than the ShapeNet or ModelNet test sets (which contain several thousands shapes).
(E5) Shape variety with defects. We evaluate both neural and traditional optimization-based methods that do not require parameterization based on a training set on Berger et al.’s synthetic range-scanning dataset.
Evaluation of all methods on a common ground. Last, we evaluate all types of methods on the same, small-enough datasets, featuring both synthetic and real point clouds.
(E6) Synthetic data with ground truth (quantitative evaluation). We compare learning- and optimization-based methods for the MVS-scanned version of five shapes from Berger et al.’s dataset, which do not correspond to any of ShapeNet’s categories but are comparable in complexity. We use the models trained in E1.
(E7) Real data without ground truth (qualitative evaluation). Finally, we compare the learning-based network trained in E1 to optimization-based methods on real point clouds obtained by MVS and range scanning. As these real shapes do not have a ground truth surface, we only perform qualitative evaluation.
5.4 Evaluated surface reconstruction methods
In this section, we offer a concise overview and some specifics regarding the implementation of the methods evaluated in our benchmark. For a more comprehensive analysis, please refer to our detailed survey presented in Sec. 4.
To ensure a fair evaluation of traditional methods, when running experiments with a training set (E1-E4), we perform a grid-search on the parameters and select the configuration with the best mean volumetric IoU on the training set. In the experiments E5-E7, we use default parameters for all optimization methods.
Learning-based methods. The parameters of these methods are learned on a training set.
ConvONet [11] first extracts point features and averages them on cells of three 2D grids (2D variant), or one 3D grid (3D variant). It then uses convolutions to capture the local geometry and predicts the occupancy of query points from interpolated grid features.
DGNN [15] uses a graph neural network to predict the occupancy of Delaunay cells.
Neural optimization-based methods. These methods are optimized for each input shape independently, and do not rely on any training set.
IGR [54] learns an implicit function for an input cloud based on its points’ position and normals. We train the model for iterations for each scan.
LIG [13] trains an autoencoder from ground truth signed distance function and only retains its decoder for inference. As the training code is unavailable, we use the available decoder pretrained on ShapeNet (without noise). The input point clouds are augmented with new points along each point’s jet-estimated normals [89], whose occupancy is determined by the normals orientation. We use the proposed post-processing to remove falsely enclosed volumes
P2M [40] iteratively moves the vertices of an initial mesh to fit a point cloud.
SAP∗ [55] is the variant of SAP that is purely optimization-based, rather than learning-based. It uses non-learned normals and densification.
Note that while LIG employs a pretrained network, it also involves optimizing a unique function for each new input point cloud. Consequently, for the purposes of our analysis and classification, we categorize LIG as an optimization-based method, acknowledging its hybrid nature.
Traditional optimization-based methods. These test-of-time methods do not rely on neural networks or training, but nonetheless have a few parameters, which can be set according to default values or adjusted depending on given ground-truth data (akin to a train set).
SPSR [27] is a classic traditional method which approximates the surface as a level-set of an implicit function estimated from point positions and normal information. We use the sensor position to orient jet-estimated normals [89]. In experiments E1-E4, we perform a grid-search on the train set with octree depths of , and boundary conditions in . In experiments E5-E7, we use an octree of depth and Dirichlet boundary condition. We also experimented with post-processing the reconstructed surface with the trimming tool provided by the authors but could not find parameters that consistently improve the evaluation metrics.
RESR [46] is a classical graph-cut-based method exploiting on visibility information. For experiments E1-E4, we perform a grid-search on the train set with point weights and , and regularization strength . In the optimization setting, we use the parametrization suggested by the authors: point weights and and regularization strength .
5.5 Evaluation metrics
We want the reconstructed surface to be as close as possible to the ground-truth surface in terms of geometry and topology. To measure this closeness we use various metrics described below.
Geometric metrics. We evaluate reconstructions with the volumetric intersection over union (IoU), symmetric Chamfer distance (CD), and normal consistency (NC):
Volumetric IoU: We denote by the space inside the ground-truth surface , and the space inside the reconstructed surface . The volumetric IoU estimates the intersection over union between and . To approximate this value, we sample a set of points in the union of the bounding boxes of and and calculate the following:
Chamfer distance: We sample a set of points and on the facets of the ground-truth mesh and the reconstructed mesh, respectively, with . We approximate the symmetric Chamfer distance between and as follows:
Normal consistency: Let be the unit normal of a point on a mesh, and the Euclidean scalar product in . The normal consistency is estimated as:
Topologic metrics. All surfaces considered in our benchmark have a single component, and are closed and manifold. In consequence, we can measure the geometric quality of the reconstructed surfaces with the following indices:
Number of components: The reconstructed surfaces should also have one component.
Number of boundary edges: All reconstructed meshes should have no boundary edges, i.e., edges that belong to only one facet.
Number of non-manifold edges: All the edges of the reconstructed meshes should be manifold.
6 Experiments
We report here the results of the series of experiments presented in Sec. 5.3 and designed to assess the robustness and performance of various deep and non-deep surface reconstruction methods from point clouds (cf. Tab. I). We start in Sec. 6.1, for methods with dataset-driven parameterization, by measuring the impact of various inconsistencies between train and test sets. We then evaluate in Sec. 6.2 the performance of optimization-based methods without dataset-driven parameterization. Finally, we compare in Sec. 6.3 all evaluated methods, both learning- and optimization-based, on synthetic and real point acquisitions.
6.1 Surface reconstruction with dataset-driven parameterization (E1 - E4)
| Volumetric IoU (%) [] | Normal consistency (%) [] | ||||||||||
| Method | E1 | E2 | E3 | E4 | Mean | E1 | E2 | E3 | E4 | Mean | |
| ConvONet2D | [11] | 85.0 | 47.3 | 79.3 | 68.3 | 70.0 | 92.7 | 76.4 | 90.0 | 87.8 | 86.7 |
| ConvONet3D | [11] | 84.8 | 15.1 | 83.6 | 51.0 | 58.6 | 93.0 | 71.8 | 93.1 | 82.5 | 85.1 |
| SAP | [55] | 88.7 | 59.8 | 89.2 | 54.9 | 73.2 | 93.5 | 86.7 | 94.1 | 87.1 | 90.4 |
| DGNN | [15] | 84.5 | 38.1 | 87.0 | 84.4 | 73.5 | 85.4 | 68.8 | 88.5 | 85.5 | 82.1 |
| POCO | [17] | 89.5 | 8.74 | 90.6 | 40.9 | 57.4 | 93.6 | 75.6 | 94.2 | 82.9 | 86.6 |
| †SPSR | [27] | 77.1 | 80.7 | 80.7 | 74.6 | 78.3 | 87.7 | 83.2 | 89.1 | 87.0 | 86.9 |
| †RESR | [46] | 80.3 | 60.4 | 83.9 | 80.3 | 76.2 | 81.0 | 73 | 84.6 | 81.0 | 79.9 |
| Chamfer distance () [] | Number of components [] | ||||||||||
| Method | E1 | E2 | E3 | E4 | Mean | E1 | E2 | E3 | E4 | Mean | |
| ConvONet2D | [11] | 0.55 | 7.51 | 1.00 | 0.98 | 2.51 | 1.6 | 34.8 | 2.6 | 3.2 | 10.6 |
| ConvONet3D | [11] | 0.55 | 10.90 | 0.76 | 2.44 | 3.66 | 1.4 | 13.6 | 1.6 | 1.5 | 4.5 |
| SAP | [55] | 0.44 | 2.09 | 0.55 | 0.92 | 1.00 | 2.7 | 86.0 | 3.5 | 10.5 | 25.7 |
| DGNN | [15] | 0.55 | 2.54 | 0.63 | 0.55 | 1.07 | 1.3 | 16.1 | 1.1 | 1.3 | 5.0 |
| POCO | [17] | 0.42 | 10.50 | 0.52 | 1.32 | 3.19 | 2.3 | 178 | 2.8 | 16.3 | 49.9 |
| †SPSR | [27] | 0.80 | 0.66 | 0.87 | 0.89 | 0.81 | 9.3 | 185 | 11.1 | 3.2 | 52.0 |
| †RESR | [46] | 0.67 | 6.97 | 0.75 | 0.67 | 2.27 | 1.2 | 9.0 | 1.0 | 1.2 | 3.1 |
| Number of boundary edges [] | Number of non-manifold edges [] | ||||||||||
| Method | E1 | E2 | E3 | E4 | Mean | E1 | E2 | E3 | E4 | Mean | |
| ConvONet2D | [11] | 0 | 0 | 0 | 0 | 0 | 0 | 0 | 0 | 0 | 0 |
| ConvONet3D | [11] | 0 | 0 | 0 | 0 | 0 | 0 | 0 | 0 | 0 | 0 |
| SAP | [55] | 0 | 0 | 0 | 8.4 | 2.1 | 0 | 0 | 0 | 0 | 0 |
| DGNN | [15] | 0 | 0 | 0 | 0 | 0 | 1.4 | 2.2 | 0.6 | 1.7 | 1.5 |
| POCO | [17] | 0 | 121 | 0 | 41.7 | 40.7 | 0 | 0 | 0 | 0 | 0 |
| †SPSR | [27] | 0 | 0 | 0 | 0 | 0 | 0 | 0 | 0 | 0 | 0 |
| †RESR | [46] | 0 | 0 | 0 | 0 | 0 | 9.4 | 28.5 | 8.47 | 9.4 | 13.9 |
Input |
CONet2D [11] |
CONet3D [11] |
SAP [55] |
DGNN [15] |
POCO [17] |
†SPSR [27] |
†RESR [46] |
Ground truth |
| (E1) In-distribution | (E2) Out-of-distribution: | (E3) Out-of-category: | (E4) Out-of-category: | (E6) Out-of-category: | |
| higher density, outliers | less complex shapes | more complex shapes | similar complex shapes |
We examine the precision and versatility of recent supervised learning-based methods, as well as two traditional optimization-based methods, whose parameters were trained or tuned on a training set. We report qualitative results in Fig. 7, and quantitative results in Tab. II.
[…]
E3: Out-of-distribution - unseen classes (simpler). In the task of reconstructing simpler, out-of-category shapes from ModelNet (E3), neural implicit field methods stand out for their visually superior reconstructions. SAP and POCO lead in terms of quantitative performance. The interpolating approach DGNN outperforms ConvONet in this experiment, demonstrating its generalization capabilities.
E4: Out-of-distribution - unseen classes (more complex). In this experiment, most methods overfit the simpler training shapes from ModelNet shapes and subsequently struggle with reconstructing the more complex shapes from ShapeNet. The non-learning-based method SPSR is also affected as its parameters are tuned on ModelNet. Specifically, the optimalo octree depth for ModelNet reconstructions is , as opposed to for ShapeNet, resulting in a decrease in volumetric IoU from in E1 to in E4. In contrast, the parameter tuning for RESR remains consistent across both datasets. DGNN stands out as the only learning-based method that does not exhibit overfitting on the ModelNet dataset, leading to the most successful outcomes both quantitatively and qualitatively. In fact, its performance is on par with its results when directly trained on the ShapeNet dataset. We attribute this behavior to the local nature of DGNN, which only consider small subgraphs during training and inference.
Methods Performance. We propose an analysis of the performance of each model:
ConvONet perform best when both the training and test sets share identical point cloud characteristics and shape categories. On more complicated settings, this method performs below traditional approaches.
SAP stands out for its superior reconstructions and robustness against outliers, partly due to its unique capability of explicitly predicting normals. This feature contributes to SAP achieving the highest mean normal consistency across all experiments.
DGNN, with its blend of local learning and global regularization, consistently delivers competitive results in nearly all scenarios, with the notable exception of the outlier-rich environment in E2. Overall, it achieves the highest volumetric IoU amongst learning-based methods when averaged across all experiments.
POCO’s local attention-based learning mechanism yields optimal outcomes when reconstruction does not require adapting to unseen domains. It particularly excels in experiments where the point cloud characteristics remain consistent between training and testing (E1, E3, E4). However, POCO is significantly impacted by outliers, as evidenced in E2, and shows a tendency to overfit on simpler training shapes, as observed in E4. Notably, both POCO and SAP produce watertight reconstructions, with boundary edges only appearing where the reconstruction intersects the bounding box.
[…]
spsr demonstrates a high degree of robustness to various defects and shape complexities, yielding consistently good results across different metrics. However, it tends to produce less compact reconstructions with a higher number of components.
RESR, while slightly less robust to outliers and now relatively old, achieves a mean volumetric IoU higher than any of the learning-based methods, and presents the most compact reconstructions with an average of 2.7 components only. However, this approach produces a significant number of non-manifold edges.
6.2 Surface reconstruction without dataset-driven parameterization (E5)
| Volumetric IoU (%) [] | Normal consistency (%) [] | ||||||||||||
| Method | LR | HR | HRN | HRO | HRNO | Mean | LR | HR | HRN | HRO | HRNO | Mean | |
| IGR | [54] | 80.8 | 92.5 | 83.6 | 63.7 | 62.7 | 76.7 | 88 | 96.3 | 83.9 | 77.8 | 71.5 | 83.5 |
| LIG | [13] | 46.9 | 50.3 | 63.9 | 66 | 63.8 | 58.2 | 88.7 | 92.2 | 89.0 | 77 | 75.2 | 84.4 |
| P2M | [40] | 75.2 | 83.3 | 75.5 | 71.3 | 67.8 | 74.6 | 86.3 | 92.2 | 88.1 | 84.5 | 82.1 | 86.6 |
| †SAP∗ | [55] | 75.6 | 89.1 | 72.4 | 55.3 | 34.9 | 65.4 | 83.4 | 94.8 | 61.6 | 74.5 | 55.3 | 73.9 |
| †SPSR | [27] | 77.7 | 90.2 | 82.8 | 90.3 | 82.1 | 84.6 | 88.1 | 96 | 88.1 | 96.2 | 85.8 | 90.9 |
| †RESR | [46] | 81.3 | 93.4 | 80.1 | 93.4 | 79.1 | 85.5 | 87.6 | 96 | 66.3 | 94.9 | 66.5 | 82.3 |
| Chamfer distance () [] | Number of components [] | ||||||||||||
| Method | LR | HR | HRN | HRO | HRNO | Mean | LR | HR | HRN | HRO | HRNO | Mean | |
| IGR | [54] | 0.67 | 0.32 | 0.55 | 7.96 | 7.72 | 3.45 | 6.8 | 1.2 | 35.2 | 44 | 97.4 | 36.9 |
| LIG | [13] | 0.75 | 0.58 | 0.78 | 7.89 | 7.80 | 3.56 | 1 | 1 | 1 | 1.6 | 1 | 1.12 |
| P2M | [40] | 0.82 | 0.47 | 0.73 | 1.53 | 2.13 | 1.13 | 1.2 | 1 | 1.2 | 1.4 | 1.6 | 1.28 |
| †SAP∗ | [55] | 0.85 | 0.32 | 0.70 | 3.99 | 3.93 | 1.96 | 73.2 | 85.6 | 937 | 1800 | 1.60 | 971 |
| †SPSR | [27] | 0.79 | 0.37 | 0.57 | 0.36 | 0.61 | 0.54 | 1.2 | 1.6 | 3.6 | 3.8 | 20.2 | 6.08 |
| †RESR | [46] | 0.64 | 0.31 | 0.61 | 0.34 | 0.64 | 0.51 | 1 | 1 | 1.2 | 1.2 | 1 | 1.08 |
| Number of boundary edges [] | Number of non-manifold edges [] | ||||||||||||
| Method | LR | HR | HRN | HRO | HRNO | Mean | LR | HR | HRN | HRO | HRNO | Mean | |
| IGR | [54] | 0 | 0 | 0 | 0 | 0 | 0 | 0 | 0.8 | 0.8 | 5.2 | 4.2 | 2.2 |
| LIG | [13] | 69 | 42.8 | 17.2 | 0 | 0 | 25.8 | 0 | 0 | 0 | 0 | 0 | 0 |
| P2M | [40] | 0 | 0 | 0 | 0 | 0 | 0 | 0 | 0 | 0 | 0 | 0 | 0 |
| †SAP∗ | [55] | 0 | 0 | 0 | 0 | 449 | 89.8 | 0 | 0 | 0 | 0 | 0 | 0 |
| †SPSR | [27] | 0 | 0 | 0 | 0 | 0 | 0 | 0 | 0 | 0 | 0 | 0 | 0 |
| †RESR | [46] | 0 | 0 | 0 | 0 | 0 | 0 | 1 | 5.8 | 24.4 | 3.8 | 22 | 11.4 |
Setting. We now evaluate methods that do not learn or tune their parameters on a training set. We consider neural-network-based optimization approaches that learn a new function to fit each input point cloud. We also assess the performance of non-learning-based methods, using their default parameters for a fair evaluation. These methods are tested on the benchmark shapes from Berger et al. [19], which uses synthetic range scans with varying resolutions, noise levels, and outlier proportions.
[…]
6.3 Learning- and optimization-based surface reconstruction from synthetic MVS point clouds (E6)
| Vol. IoU | Normal consist. | Chamfer dist. | Components | BE | NME | |||
| Method | (%) [] | (%) [] | () [] | [] | [] | [] | ||
| Learning | ConvONet2D | [11] | 65.1 | 78.0 | 1.43 | 3.6 | 0 | 0 |
| ConvONet3D | [11] | 76.4 | 87.2 | 0.89 | 2.6 | 0 | 0 | |
| SAP | [55] | 78.3 | 89.0 | 0.73 | 5.6 | 0 | 0 | |
| DGNN | [15] | 82.9 | 85.2 | 0.59 | 1 | 0 | 0.4 | |
| POCO | [17] | 83.9 | 89.5 | 0.58 | 2 | 0 | 0 | |
| Optimization | IGR | [54] | 78.3 | 83.8 | 0.78 | 15.4 | 0 | 0.4 |
| LIG | [13] | 45.7 | 86.6 | 0.83 | 1 | 65.6 | 0 | |
| P2M | [40] | 74.5 | 85.0 | 0.77 | 2 | 0 | 0 | |
| †SAP | [55] | 71.9 | 77.0 | 0.81 | 133 | 0 | 0 | |
| †SPSR | [27] | 77.6 | 86.4 | 0.79 | 8 | 0 | 0 | |
| †RESR | [46] | 79.4 | 80.8 | 0.67 | 1 | 0 | 9.6 | |
In this other experiment, we compare learning- and optimization-based reconstruction methods on the same dataset. We scan the shapes of Berger et al. [19] with the same synthetic scanning method as in E5. We consider learning-based models trained on shapes from ShapeNet scanned with the same procedure, network-based optimization models without training, and traditional optimization approaches. We report the results in Tab. IV and provide visualizations in the supplementary material. Among the learning-based methods, DGNN and POCO achieve the most accurate surface reconstructions. Consistent with findings from E5, RESR emerges as the most effective among the optimization-based techniques. However, learning-based models achieve higher precision by leveraging priors from ShapeNet thanks to the similar scanning protocol.
6.4 Learning- and optimization-based surface reconstruction from real point clouds (E7)
Learning |
Optimization |
Learning |
Optimization |
Learning |
Optimization |
In this final experiment, we qualitatively assess the performance of learning- and optimization-based methods on real MVS and range-scanned point clouds (without ground truth). For the learning-based approaches, we use models trained on synthetic MVS scans from ShapeNet (E4), while optimization-based methods are applied independently to each point cloud. Our findings are presented in Fig. 8.
[…]
This experiment reinforces our findings from synthetic point clouds: learning-based methods are vulnerable to unfamiliar defects, whereas traditional methods offer a robust and reliable baseline.
6.5 Runtimes
The detailed runtimes for the methods evaluated in the learning-based experiments are presented in Tab. V. SAP stands out as the fastest reconstruction method. DGNN also exhibits efficient runtimes, whereas POCO tends to be slower, attributable primarily to its intensive and costly need to search for query neighbors. Additionally, we include runtime comparisons for P2S. Due to its extensive computational demands during both training and inference, P2S was excluded from experiments E1 through E4.
6.6 Summary and analysis
When the point clouds in the training and test sets have similar characteristics, learning-based methods produce high-quality reconstructions and even show robustness to noise and missing data. They are able to generalize to unseen shape categories, provided that the training set is sufficiently large (30k shapes in our experiments) and contains shapes of sufficient complexity. This ability implies that the methods primarily learn priors associated with point cloud attributes rather than the shapes themselves. However, learning-based methods do not produce satisfying results when the training shapes are too simple, or when the point clouds include unknown defects, such as outliers. Hybrid approaches such as SAP or DGNN achieve higher robustness to domain shifts and shorter reconstruction times. With the exception of IGR, the newer optimization-based methods we tested do not exhibit strong resilience to acquisition defects and generally do not outperform the traditional methods SPSR and RESR.
7 Conclusion
Surface reconstruction from point clouds is a central subject in digital geometry processing, continually evolving with great strides in acquisition technologies and innovative approaches to surface reconstruction and analysis. In this paper, we have conducted an extensive survey of the field, benchmarking across various datasets a range of recent learning-based and optimization-based methods, alongside established traditional techniques. Our findings reveal that learning-based methods demonstrate compelling performance when tested on point clouds bearing resemblance to their training sets. However, their effectiveness is contingent upon training with shapes of comparable complexity to the test set and they exhibit limited robustness to out-of-distribution acquisition defects. While they may lag behind in in-distribution scenarios, traditional methods consistently offer robust and reliable performance across a broader range of conditions. We also observe that recent learning-free optimization-based methods are outperformed by these traditional methods in almost all settings.
Real-world scenes often include a variety of complex acquisition defects. Consequently, learning-based methods face challenges in reconstructing such complex point clouds accurately. Thus, the creation or compilation of training data that encompasses a wide array of complex shapes, scanned with realistic defects, becomes crucial yet poses significant challenges. Future directions in learning-based surface reconstruction should prioritize training on point clouds with realistic acquisition defects, such as those common in typical sensor and acquisition setups, or focus on enhancing the resilience of these methods to novel and unseen defects.